OpenStudy (sweetdanny):

I have this question in vector calculus. pls help I tried everything I know....

OpenStudy (sweetdanny):

pls check attachment: To find surface integral ????

OpenStudy (sweetdanny):

I will give a medal and best feedback :)

OpenStudy (kainui):

Try drawing a picture of the lines \(y=3-z^2\) , \(x=0\) and \(z=5\)

OpenStudy (sweetdanny):

inverted parabola on yz plane pushed 3units up

OpenStudy (sweetdanny):

I used double integral(F.ds) but boundaries for x??? I cant tell

OpenStudy (kainui):

Strange. It seems like there's no lower bound on z and no upper bound on x. But the way they've written it is weird, I wonder if they're just lazy. If you look carefully they have NOT written: \[0 \le x\] and \[z\le 5\] It looks like they've confusingly written \[0 \le x, z \le 5\] Is that supposed to mean \[0 \le x\le 5\] and\[0 \le z \le 5\]? That would definitely fix the problem with there not being any bounds but seriously I have no idea why someone would do something so idiotic in writing a question up like this... It's confusing enough as it is. Ugh

OpenStudy (kainui):

The other alternative is just to take it to mean it's an unbounded integral so it's like an infinite sheet perhaps.

OpenStudy (kainui):